Q2.(A) Given the following functions

           Q=10-2P2                               Q= 5+3P

Find

i)                   Nature of each function

ii)                 Market price and Market quantity

iii)               Elasticity of Qd and Qs at equilibrium price



1
Expert's answer
2021-04-07T10:41:58-0400
"Q_D=10-2p^2"

"Q_S=5+3p"

"10-2p^2=5+3p"

"p=1"

"Q=8"

"E_D=-\frac {1}{2}"

"E_S=\frac {3}{8}"